OpenStudy (anonymous):

Graph the circle (x^2)+(y^2)-8x+6y+21=0

OpenStudy (anonymous):

Ok.. How are you with completing squares?

OpenStudy (anonymous):

medal me.!!!

OpenStudy (anonymous):

To graph it by hand, however, you would need to complete the squares... Look at the variables to the first power, like x and y, and use the constants to help you complete the square.

OpenStudy (anonymous):

add numbers to both sides of the equations as you do it, so that you can balancing it out. The end result would be like (x-a)^2 + (y-b)^2 = r^2, where r is the radios, and (a,b) is the center of the circle.

OpenStudy (anonymous):

Im not that good with completing squares. I don't remember the process

OpenStudy (anonymous):

Well.. You need to divide the constant by 2, and then square that. So if the equation is like x^2 - 8x, you divide the first power x by 2, to get -4, and then you square that number, which will give you 16. Add 16 to both sides of the equation. and you will get x^2 -8x +16 = 16. You now can complete the square on the left side, which will be (x-4)^2 = 16.